OpenStudy (anonymous):

HELP PLEASSSEE!! Factorize and solve for X: 20+9x=x^2=0

OpenStudy (anonymous):

too many equal signs

OpenStudy (anonymous):

oh sorry! wait

OpenStudy (anonymous):

20+9x+x^2=0

OpenStudy (anonymous):

\[(x+5)(x+4)=0\]

hero (hero):

x^2 + 9x + 20 = 0 x^2 + 5x + 4x + 20 = 0 x(x+5)+4(x+5) = 0 (x+5)(x+4) = 0

OpenStudy (precal):

x=-5 and x=-4